Grade 12th passAlgebra

the value of cos 20+8 sin70*sin 50*sin 10/sin^2(80) is equal to

twisha
8 Years agoGrade 12th pass

1 Answer

Arun
8 Years ago
Dear student
 
Sin10sin50sin70 = 2sin10cos10sin50sin70/2cos10
 
= sin20sin50sin70/2cos10 = 2sin20cos20cos40/4cos10 
 
= 2sin40cos40/8cos10 = sin80/8sin80 = 1/8 
 
cos20 + 8sin10sin50sin70 = cos20 + 1 = 2cos²10 = 2sin²80 
 
(cos20 + 8sin70sin50sin10)/sin²80 = 2.
 
 
 
Regards
Arun (askIITians forum expert)